OpenStudy (anonymous):

Solve the system of equations using the substitution method. what is x and y 2x + y = 8 y = -x+5

OpenStudy (anonymous):

answer choices ?

OpenStudy (anonymous):

there aren't any

OpenStudy (anonymous):

ok Ill do the best I can :D

OpenStudy (anonymous):

ok thanks ! : P

OpenStudy (anonymous):

meet me here http://www.twiddla.com/1248793

OpenStudy (mertsj):

Replace y in the first equation with -x+5 from the second equation.

HanAkoSolo (jamierox4ev3r):

it is explicitly stated that y=-x+5. Therefore, you can plug this in to the other equation to create a big equation. Let me show you what i mean

HanAkoSolo (jamierox4ev3r):

2x+(-x+5)=8

OpenStudy (anonymous):

ok i got x=4 and y=1 is that right ?

HanAkoSolo (jamierox4ev3r):

2x-x+5=8 x+5=8 -5 -5 x=3

HanAkoSolo (jamierox4ev3r):

well, so far i got that x=3

HanAkoSolo (jamierox4ev3r):

rlly all i did was what @Mertsj did

HanAkoSolo (jamierox4ev3r):

then you find the y-variable by plugging in the x value into both equations

OpenStudy (anonymous):

ok if you dont mind can you draw it on the drawing bored or in http://www.twiddla.com/

HanAkoSolo (jamierox4ev3r):

y=-3+5 y=2 2(3)+y=8 6+y=8 -6 -6 y=2

HanAkoSolo (jamierox4ev3r):

by plugging the x-value into both equations, we have established that y=2 and x=3

HanAkoSolo (jamierox4ev3r):

but sure i will do it on the drawing board if u want

OpenStudy (anonymous):

ok thanks

HanAkoSolo (jamierox4ev3r):

hmm it doesn't work on my laptop can i just draw it on here?

HanAkoSolo (jamierox4ev3r):

@ineedhelp15 ?

OpenStudy (anonymous):

ok

HanAkoSolo (jamierox4ev3r):



HanAkoSolo (jamierox4ev3r):

ok phew that took a while. does that make sense? @ineedhelp15 ?

OpenStudy (anonymous):

yes :) thanks !

HanAkoSolo (jamierox4ev3r):

haha np nice working wit you :)